OpenStudy (anonymous):

In a word game, you choose a tile from a bag, replace it, and then choose another. If there are 21 vowels and 15 consonants, what is the probability you will choose a consonant and then a vowel?

OpenStudy (anonymous):

i got 3/144

OpenStudy (anonymous):

Total constants = 15 Total vowels = 21 Total letters = 15 + 21 = 36 Probability of a consonant on first place = 15/36 Probability of a vowel on second place = 21/36 Probability of getting a consonant and then a vowel (or vice-versa): (15/36)(21/36) = 315/1296 = 35/144 can you check it plzz @shubhamsrg

OpenStudy (shubhamsrg):

absolutely correct,35/144 is the correct ans

OpenStudy (anonymous):

thnk